# Time-dependent Predictive Accuracy Metrics in the Context of Interval Censoring and Competing Risks

### v1

@@ Abstract
Evaluating the performance of a prediction model is a common task in medical statistics. Standard accuracy metrics require the observation of the true outcomes. This is typically not possible in the setting with time-to-event outcomes due to censoring. Interval censoring, the presence of time-varying covariates, and competing risks present additional challenges in obtaining those accuracy metrics. In this study, we propose two methods to deal with interval censoring in a time-varying competing risk setting: a model-based approach and the inverse probability of censoring weighting (IPCW) approach, focusing on three key time-dependent metrics: area under the receiver-operating characteristic curve (AUC), Brier score, and expected predictive cross-entropy (EPCE). The evaluation is conducted over a medically relevant time interval of interest, $[t,\Delta t)$. The model-based approach includes all subjects in the risk set, using their predicted risks to contribute to the accuracy metrics. In contrast, the IPCW approach only considers the subset of subjects who are known to be event-free or experience the event within the interval of interest. we performed a simulation study to compare the performance of the two approaches with regard to the three metrics.
@@ 1 Introduction
In settings with interval censoring, the situation is even more complicated because also for patients who are known to have experienced the event of interest during the follow-up, the exact event time is unknown.(Zhang and Sun, 2010) Consequently, the ambiguity of the underlying unobserved event times bring challenges in defining cases (patients who experience the event before a timepoint or within a particular time interval) and controls (patients who “survived” up to a timepoint or time interval). The identification of cases and controls is essential in formulating the sensitivity and specificity of the prediction model, based on which the ROC curve is calculated. Tsouprou (2015) described a method to order the event times of any two interval-censored subjects via a joint density of the two estimated event times when calculating the concordance index. Time-varying covariates have been drawing greater attention recently in medical applications. The evaluation then needs to be done for a specific time point or a time interval, and the metrics should be a function of time. To accommodate this, time-dependent evaluation metrics were developed, which allow for the dynamic determination of cases and controls.(Bansal and Heagerty, 2018) Researchers can then evaluate the model using its predicted risk for an event before a time $t+\Delta t$ conditional on covariate information and being event-free up until a time $t$. This can be seen as the risk of experiencing an event in a clinically relevant interval $[t,t+dt)$. Last but not least, another common complication is the presence of competing events: patients who are observed with the competing event might still have experienced the (yet unobserved) interval-censored event. This has to be taken into account when obtaining a risk estimate but also further complicates the determination of whether the patient is a case or control during model evaluation. Although others have proposed adaptions of the definition of sensitivity being event-specific,(Saha and Heagerty, 2010; Dey et al., 2020) the combination of the abovementioned difficulties brings forth challenges in the model evaluation.
@@
This work aims to extend the formulation of time-dependent predictive accuracy metrics, specifically the AUC, Brier score and EPCE, in the context of interval-censored data with competing risks. We present the adapted accuracy metrics using the Canary Prostate Active Surveillance Study (PASS).(Newcomb et al., 2016) The primary event of interest is prostate cancer progression, which is interval-censored between a negative biopsy and a positive biopsy. Since the start of treatment means the end of active surveillance, cancer progression while on active surveillance can no longer be observed, thus the treatment initiation constitutes a competing risk. The example prediction model, based on the motivating data, is a joint model (ICJM) for repeatedly measured prostate-specific antigen (PSA) and the time to both the primary and competing events.(Yang et al., 2024a) We explore and extend two methods to evaluate the predictive performance in the context of interval censoring and competing risk: a model-based approach and the IPCW approach. The former takes advantage of all patients at risk at the beginning of an interval of interest, $[t,t+dt)$, to determine the probabilities of being a case or control for the particular interval and weighs those patients based on their predicted risks from the model or algorithm itself. The latter uses only the subset of the sample for which it is known whether they are a case or a control, and weights this subset to represent the whole sample. We compare the AUC and Brier scores using both approaches and explore the impact of model misspecification as well as different patterns of interval censoring. For the model-based approach, we also evaluate the EPCE. The evaluation metrics are applicable to general statistical models as well as other prediction algorithms such as machine learning techniques.
@@ 2.1 Notation
Motivated by the Canary PASS data, we denote $\delta_{i}$ as the observed event type of patient $i$, with $\delta_{i}=1$ being cancer progression, $\delta_{i}=2$ early treatment, and $\delta_{i}=3$ right censoring (e.g., due to loss to follow-up or at the time of data extraction). We distinguish between the true (but unobserved) time of cancer progression ($T_{i}^{\textsc{prg}*}$), the true time of early treatment initiation ($T_{i}^{\textsc{trt}*}$), and the censoring time ($T^{\textsc{cen}}_{i}$). For patients recorded with the interval-censored event, cancer progression ($T_{i}^{\textsc{prg}*}<\min\left\{T^{\textsc{cen}}_{i},T_{i}^{\textsc{trt}*}\right\}$), it is only known that progression happened between the last progression-free biopsy (i.e., the observed progression-free time), $T^{\textsc{prg-}}_{i}$, and the biopsy time at which progression was detected, $T^{\textsc{prg+}}_{i}$. For the patients who started early treatment ($T_{i}^{\textsc{trt}*}<\min\left\{T^{\textsc{cen}}_{i},T_{i}^{\textsc{prg+}}\right\}$), progression still might have occurred in the interval between the last progression-free biopsy and the treatment initiation time $T_{i}^{\textsc{trt}}=T_{i}^{\textsc{trt}*}$ but was not detected. For patients who are censored at $T^{\textsc{cen}}_{i}$, the corresponding progression-risk interval is between the last progression-free biopsy and infinity.
We denote the vector of repeated observations of the PSA level with $\boldsymbol{\mathcal{Y}}_{i}$ for the $i$-th patient. We distinguish the training dataset, used to fit the prediction model, and the test dataset, in which the prediction model is being evaluated. The training set is denoted by $\boldsymbol{\mathcal{D}}_{n}$ and includes the longitudinal PSA measurements, event times and other (baseline) covariates of $n$ subjects. We also define an interval of interest for prediction, $[t,t+\Delta t)$. The quantity of interest is the progression-specific risk at $t+\Delta t$ conditional on no event occurring before time $t$, denoted by $\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)$ and can be formulated as
$\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)=\Pr\Bigg{[}T^{\textsc{prg}*}_{i}\leq t+\Delta t,T^{\textsc{prg}*}_{i}<T^{\textsc{trt}*}_{i}\mid T^{\textsc{prg}*}_{i}>t,T^{\textsc{trt}*}_{i}>t,\boldsymbol{\mathcal{X}}_{i}(t),\boldsymbol{\mathcal{D}}_{n}\Bigg{]},$
where $\boldsymbol{\mathcal{X}}_{i}(t)$ includes the longitudinal PSA measurements up until time $t$ and all other (baseline) covariates used in the prediction. Typically, a subject is considered as a case if their event occurs within the interval of interest, and as a control if the event is after that interval. In the test set, the $n_{t}$ patients who have not progressed, been treated, or censored before time $t$ form the risk set and are considered in calculating the evaluation metrics in the following subsections.
@@ 2.3 Sensitivity, specificity and AUC
$\text{sen}^{\textsc{prg}}(t,\Delta t,c)$ $=\Pr\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\geq c\mid T^{\textsc{prg*}}_{i}\geq t,T^{\textsc{prg*}}_{i}<t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}$ (1)
$=\frac{\Pr\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)\geq c,T^{\textsc{prg*}}_{i}\geq t,T^{\textsc{prg*}}_{i}<t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}}{\Pr\{T^{\textsc{prg*}}_{i}\geq t,T^{\textsc{prg*}}_{i}<t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}},$
where $c$ is a value in $[0,1]$. Analogously, the progression-specific specificity, used to evaluate the ability to identify the negative instances, is defined as
$\text{spe}^{\textsc{prg}}(t,\Delta t,c)$ $=\Pr\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\mid T^{\textsc{prg*}}_{i}\geq t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}$ (2)
$=\frac{\Pr\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c,T^{\textsc{prg*}}_{i}\geq t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}}{\Pr\{T^{\textsc{prg*}}_{i}\geq t+\Delta t,T^{\textsc{prg*}}_{i}<T^{\textsc{trt*}}_{i}\}},$
@@
In the model-based approach, all subjects at risk at time $t$ are considered when calculating the sensitivity, but they contribute with different weights, $\mathcal{W}^{\textsc{m}}_{i}$, that depend on their estimated probability of experiencing cancer progression during the interval of interest. The weights are derived from the cumulative incidence function estimated by the model, and are, thus, named model-based weights. The estimated model-based progression-specific sensitivity can be written as
@@
$\mathcal{W}^{\textsc{m}}_{i}=\begin{cases}\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{prg+}}_{i}>t,T^{\textsc{prg+}}_{i}<t+\Delta t\quad\scriptstyle{(1a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{trt}}_{i}>t,T^{\textsc{trt}}_{i}<t+\Delta t\quad\scriptstyle{(1b)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(1c)},\\ 1&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg+}}_{i}<t+\Delta t\quad\scriptstyle{(3a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{trt}}_{i}<t+\Delta t\quad\scriptstyle{(3b)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(3c)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{(2)}_{i}>t+\Delta t\quad\scriptstyle{(2a,2b,2c)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{(2)}_{i}>t+\Delta t\quad\scriptstyle{(5a,5b,5c)},\\ 0&\text{otherwise}\quad\scriptstyle{(4a,4b,4c,\text{ and the greyed-out scenarios})},\end{cases}$
where $T^{(2)}_{i}$ is $T^{\textsc{cen}}_{i}$, $T^{\textsc{prg+}}_{i}$ or $T^{\textsc{trt}}_{i}$. Likewise, the progression-specific specificity can be estimated correspondingly, by using the model-based weights (visualized in Figure 2(b))
$\widehat{\text{spe}}^{\textsc{prg}}_{\textsc{model}}(t,\Delta t,c)$ $=\frac{\sum_{i}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\}\times\mathcal{W^{\prime}}^{\textsc{m}}_{i}}{\sum_{i}\mathcal{W^{\prime}}^{\textsc{m}}_{i}},$

### v2

@@ Abstract
Evaluating the performance of a prediction model is a common task in medical statistics. Standard accuracy metrics require the observation of the true outcomes. This is typically not possible in the setting with time-to-event outcomes due to censoring. Interval censoring, the presence of time-varying covariates, and competing risks present additional challenges in obtaining those accuracy metrics. In this study, we propose two methods to deal with interval censoring in a time-varying competing risk setting: a model-based approach and the inverse probability of censoring weighting (IPCW) approach, focusing on three key time-dependent metrics: area under the receiver-operating characteristic curve (AUC), Brier score, and expected predictive cross-entropy (EPCE). The evaluation is conducted over a medically relevant time interval of interest, $[t,\Delta t)$. The model-based approach includes all subjects in the risk set, using their predicted risks to contribute to the accuracy metrics. In contrast, the IPCW approach only considers the subset of subjects who are known to be event-free or experience the event within the interval of interest. We performed a simulation study to compare the performance of the two approaches with regard to the three metrics. Furthermore, we demonstrated the three metrics using the two approaches on an example prostate cancer surveillance cohort. Risk predictions were generated from a joint model handling the interval-censored cancer progression and the competing event, early treatment, and repeatedly measured biomarkers.
@@ 1 Introduction
In settings with interval censoring, the situation is even more complicated because also for patients who are known to have experienced the event of interest during the follow-up, the exact event time is unknown.(Zhang and Sun, 2010) Consequently, the ambiguity of the underlying unobserved event times bring challenges in defining cases (patients who experience the event before a timepoint or within a particular time interval) and controls (patients who “survived” up to a timepoint or time interval). The identification of cases and controls is essential in formulating the sensitivity and specificity of the prediction model, based on which the ROC curve is calculated. Tsouprou (2015) described a method to order the event times of any two interval-censored subjects via a joint density of the two estimated event times when calculating the concordance index. Time-varying covariates have been drawing greater attention recently in medical applications. The evaluation then needs to be done for a specific time point or a time interval, and the metrics should be a function of time. To accommodate this, time-dependent evaluation metrics were developed, which allow for the dynamic determination of cases and controls.(Bansal and Heagerty, 2018) Researchers can then evaluate the model using its predicted risk for an event before a time $t+\Delta t$ conditional on covariate information and being event-free up until a time $t$. This can be seen as the risk of experiencing an event in a clinically relevant interval $[t,t+dt)$. Last but not least, another common complication is the presence of competing events: patients who are observed with the competing event might still have experienced the (yet unobserved) interval-censored event. This combination of competing risk and interval censoring is not a rare occurrence in clinical research, for instance, in oncology. Cancer progression or diagnosis is typically not directly observable and unless all-cause mortality is the event of interest, death poses a competing risk. This has to be taken into account when obtaining a risk estimate but also further complicates the determination of whether the patient is a case or control during model evaluation. Although others have proposed adaptions of the definition of sensitivity being event-specific,(Saha and Heagerty, 2010; Dey et al., 2020) the combination of the abovementioned difficulties brings forth challenges in the model evaluation.
@@
This work aims to extend the formulation of time-dependent predictive accuracy metrics, specifically the AUC, Brier score and EPCE, in the context of interval-censored data with competing risks. We present the adapted accuracy metrics using the Canary Prostate Active Surveillance Study (PASS).(Newcomb et al., 2016) Patients with low-risk prostate cancer were admitted to active surveillance instead of receiving immediate and typically invasive treatment, and closely monitored by regular biopsies and biomarker measurements (prostate-specific antigen, PSA). Only when a progression of the cancer is diagnosed, treatment is recommended for these patients. The primary event of interest is prostate cancer progression, which is interval-censored between a negative biopsy and a positive biopsy. Since the start of treatment means the end of active surveillance, cancer progression while on active surveillance can no longer be observed, thus the treatment initiation constitutes a competing risk. The example prediction model, based on the motivating data, is a joint model (ICJM) for repeatedly measured PSA and the time to both the primary and competing events.(Yang et al., 2025) We explore and extend two methods to evaluate the predictive performance in the context of interval censoring and competing risk: a model-based approach and the IPCW approach. The former takes advantage of all patients at risk at the beginning of an interval of interest, $[t,t+dt)$, to determine the probabilities of being a case or control for the particular interval and weighs those patients based on their predicted risks from the model or algorithm itself. The latter uses only the subset of the sample for which it is known whether they are a case or a control, and weights this subset to represent the whole sample. We compare the AUC and Brier scores using both approaches and explore the impact of model misspecification as well as different patterns of interval censoring. For the model-based approach, we also evaluate the EPCE. The corresponding source codes can be found on GitHub (https://github.com/ZhenweiYang96/Predicitve-accuracy-metrics). The evaluation metrics are applicable to general statistical models as well as other prediction algorithms such as machine learning techniques.
@@ 2.1 Notation
Motivated by the Canary PASS data, we denote $\delta$ as the observed event type of patients in the test set, with $\delta=1$ being cancer progression, $\delta=2$ early treatment, and $\delta=3$ right censoring (e.g., due to loss to follow-up or at the time of data extraction). We distinguish between the true (but unobserved) time of cancer progression ($T^{\textsc{prg}*}$), the true time of early treatment initiation ($T^{\textsc{trt}*}$), and the censoring time ($T^{\textsc{cen}}$). For patients recorded with the interval-censored event, cancer progression ($T^{\textsc{prg}*}<\min\left\{T^{\textsc{cen}},T^{\textsc{trt}*}\right\}$), it is only known that progression happened between the last progression-free biopsy (i.e., the observed progression-free time), $-T^{\textsc{prg-}}$, and the biopsy time at which progression was detected, $T^{\textsc{prg+}}$. For the patients who started early treatment ($T^{\textsc{trt}*}<\min\left\{T^{\textsc{cen}},T^{\textsc{prg+}}\right\}$), progression still might have occurred in the interval between the last progression-free biopsy and the treatment initiation time $T^{\textsc{trt}}=T^{\textsc{trt}*}$ but was not detected. For patients who are censored at $T^{\textsc{cen}}$, the corresponding progression-risk interval is between the last progression-free biopsy and infinity.
We denote the vector of repeated observations of the PSA level with $\boldsymbol{\mathcal{Y}}$. We distinguish the training dataset, used to fit the prediction model, and the test dataset, in which the prediction model is being evaluated. The training set is denoted by $\boldsymbol{\mathcal{D}}_{n}$ and includes the longitudinal PSA measurements, event times and other (baseline) covariates of $n$ subjects. We also define an interval of interest for prediction, $[t,t+\Delta t)$. The quantity of interest is the progression-specific risk at $t+\Delta t$ conditional on no event occurring before time $t$, denoted by $\Pi^{\textsc{prg}}(t+\Delta t\mid t)$ and can be formulated as
$\Pi^{\textsc{prg}}(t+\Delta t\mid t)=\Pr\Bigg{[}T^{\textsc{prg}*}\leq t+\Delta t,T^{\textsc{prg}*}<T^{\textsc{trt}*}\mid T^{\textsc{prg}*}>t,T^{\textsc{trt}*}>t,\boldsymbol{\mathcal{X}}(t),\boldsymbol{\mathcal{D}}_{n}\Bigg{]},$
where $\boldsymbol{\mathcal{X}}(t)$ denotes the data of the test set, including the longitudinal PSA measurements up until time $t$ and all other (baseline) covariates used in the prediction. Typically, a subject is considered as a case if their event occurs within the interval of interest, and as a control if the event is after that interval. In the test set, the $n_{t}$ patients who have not progressed, been treated, or censored before time $t$ form the risk set and are considered in calculating the evaluation metrics in the following subsections.
@@ 2.3 Sensitivity, specificity and AUC
$\text{sen}^{\textsc{prg}}(t,\Delta t,c)$ $*\displaystyle=\Pr\{\Pi^{\textsc{prg}}(t+\Delta t\mid t)\geq c\mid T^{\textsc{prg*}}\geq t,T^{\textsc{prg*}}<t+\Delta t,T^{\textsc{prg*}}<T^{\textsc{trt*}}\}$ (1)
$*\displaystyle=\frac{\Pr\{\Pi^{\textsc{prg}}(t+\Delta t\mid t)\geq c,T^{\textsc{prg*}}\geq t,T^{\textsc{prg*}}<t+\Delta t,T^{\textsc{prg*}}<T^{\textsc{trt*}}\}}{\Pr\{T^{\textsc{prg*}}\geq t,T^{\textsc{prg*}}<t+\Delta t,T^{\textsc{prg*}}<T^{\textsc{trt*}}\}},$
where $c$ is a value in $[0,1]$. Similar definitions can be found in Blanche et al. (2013, 2014). Analogously, following Equation (1) in Blanche et al. (2013), the specificity, used to evaluate the ability to identify the negative instances, is defined as
$\text{spe}(t,\Delta t,c)$ $=\Pr\{\Pi^{\textsc{prg}}(t+\Delta t\mid t)<c\mid\min(T^{\textsc{prg*}},T^{\textsc{trt*}})\geq t+\Delta t\}$ (2)
$=\frac{\Pr\{\Pi^{\textsc{prg}}(t+\Delta t\mid t)<c,\min(T^{\textsc{prg*}},T^{\textsc{trt*}}_{i})\geq t+\Delta t\}}{\Pr\{\min(T^{\textsc{prg*}},T^{\textsc{trt*}})\geq t+\Delta t\}},$
@@
In the model-based approach, all subjects at risk at time $t$ are considered when calculating the sensitivity, but they contribute with different weights, $\mathcal{W}^{\textsc{m}}_{i}$, that depend on their estimated probability of experiencing cancer progression during the interval of interest for patient $i$ in the test set. The weights are derived from the cumulative incidence function estimated by the model, and are, thus, named model-based weights. The estimated model-based progression-specific sensitivity can be written as
@@
$\mathcal{W}^{\textsc{m}}_{i}=\begin{cases}\frac{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})}{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})}&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{prg+}}_{i}>t,T^{\textsc{prg+}}_{i}<t+\Delta t\quad\scriptstyle{(1a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{trt}}_{i}>t,T^{\textsc{trt}}_{i}<t+\Delta t\quad\scriptstyle{(1b)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(1c)},\\ \frac{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})}{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})}&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(2a)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{trt}}_{i}>t+\Delta t\quad\scriptstyle{(2b)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg-}}_{i}<t+\Delta t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(2c)},\\ 1&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{prg+}}_{i}<t+\Delta t\quad\scriptstyle{(3a)},\\ \Pi^{\textsc{prg}}_{i}(T^{\textsc{trt}}_{i}\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{trt}}_{i}<t+\Delta t\quad\scriptstyle{(3b)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}>t,T^{\textsc{cen}}_{i}<t+\Delta t\quad\scriptstyle{(3c)},\\ \frac{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})}{\Pi^{\textsc{prg}}_{i}(T^{\textsc{prg+}}_{i}\mid T^{\textsc{prg-}}_{i})}&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{prg+}}_{i}>t+\Delta t\quad\scriptstyle{(5a)},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{trt}}_{i}>t+\Delta t\quad\scriptstyle{(5b},\\ \Pi^{\textsc{prg}}_{i}(t+\Delta t\mid T^{\textsc{prg-}}_{i})-\Pi^{\textsc{prg}}_{i}(t\mid T^{\textsc{prg-}}_{i})&\text{if }T^{\textsc{prg-}}_{i}<t,T^{\textsc{cen}}_{i}>t+\Delta t\quad\scriptstyle{(5c)}.\\ 0&\text{otherwise}\quad\scriptstyle{(4a,4b,4c,\text{ and the grayed-out scenarios})},\end{cases}$
It is noted that for patient groups 1a, 2a and 5a, since their progression has been detected and must be between the last negative biopsy time $T^{\textsc{prg-}}_{i}$ and the positive biopsy time $T^{\textsc{prg+}}_{i}$, the weights of being cases are the predicted progression-specific risk differences within the interval of interest rescaled relative to the corresponding risk differences between the last negative biopsy and the positive biopsy. Likewise, the specificity can be estimated correspondingly, by using the model-based weights (visualized in Figure 2(b))
$\widehat{\text{spe}}_{\textsc{model}}(t,\Delta t,c)$ $=\frac{\sum_{i}I\{\Pi^{\textsc{prg}}_{i}(t+\Delta t\mid t)<c\}\times\mathcal{W^{\prime}}^{\textsc{m}}_{i}}{\sum_{i}\mathcal{W^{\prime}}^{\textsc{m}}_{i}},$